# Linear cost and exponentially convergent approximation of Gaussian Matérn processes

### v1

@@ 2 Exponentially convergent rational approximation
where $k,r_{i}>0$ and $p_{i}<0$ for $i=1,\ldots,m$ (Bolin et al., 2024b). Let $r_{m}^{\alpha}(\cdot,\cdot)$ be the covariance function of the rational approximation induced by this approximated spectral density. Also, for simplicity, let $\varrho_{m}^{\alpha}(\cdot)$ be defined as $\varrho_{m}^{\alpha}(t-s)=r_{m}^{\alpha}(t,s)$. Based on (3), we obtain the following explicit expression of the approximated covariance function $\varrho_{m}^{\alpha}(\cdot)$.
Proposition 2.1.
@@
$\varrho_{m}^{\alpha}(h)=\varrho_{m,0}^{\alpha}(h)+\displaystyle\sum_{i=1}^{m}\varrho_{m,i}^{\alpha}(h),$ (4)
@@
$\varrho_{m,0}^{\alpha}(h)=k\sigma^{2}\cdot\begin{cases}\frac{c_{\alpha}\sqrt{4\pi}}{\kappa}1_{[h=0]}&0<\alpha<1,\\ \varrho\left(h;\lfloor\alpha\rfloor-\frac{1}{2},\kappa,\frac{c_{\alpha}}{c_{\lfloor\alpha\rfloor}}\right)&\alpha\geq 1,\end{cases}$
@@
$\varrho_{m,i}^{\alpha}(h)=r_{i}\sigma^{2}\cdot\begin{cases}\varrho\left(h;\frac{1}{2},\kappa_{i},\frac{c_{\alpha}\sqrt{\pi}}{\sqrt{1-p_{i}}}\right)&0<\alpha<1,\\ \frac{1}{p_{i}^{\lfloor\alpha\rfloor}}\varrho\left(h;\frac{1}{2},\kappa_{i},\frac{c_{\alpha}\sqrt{\pi}}{\sqrt{1-p_{i}}}\right)-\displaystyle\sum_{j=1}^{\lfloor\alpha\rfloor}\frac{1}{p_{i}^{\lfloor\alpha\rfloor+1-j}}\varrho\left(h;j-\frac{1}{2},\kappa,\frac{c_{\alpha}}{c_{j}}\right)&\alpha\geq 1.\end{cases}$
@@
[…]
Theorem 2.2.
If $\alpha>1/2$, then
$\|r_{m}^{\alpha}-r\|_{L_{2}(I\times I)}\leq C_{1}\mathbb{I}_{\alpha\notin\mathbb{N}}e^{-2\pi\sqrt{\{\alpha\}m}},$ (5)
where $r$ is given by (1) with $\nu=\alpha-\nicefrac{{1}}{{2}}$. Further, if $\alpha>1$, then
$\|r_{m}^{\alpha}-r\|_{L_{\infty}(I\times I)}\leq C_{2}\mathbb{I}_{\alpha\notin\mathbb{N}}e^{-2\pi\sqrt{\{\alpha\}m}},$ (6)
where $C_{1},C_{2}\in(0,\infty)$ are constants independent of $m$.
3 Linear cost inference
Since the spectral density in (3) is a sum of valid spectral densities $f_{m,0}^{\alpha}$ and $f_{m,i}^{\alpha}$ for $i=1,\ldots,m$, a Gaussian process with the spectral density given by (3) can be expressed as a sum of independent Gaussian processes
@@ 3 Linear cost inference
where $u_{0}$ has spectral density $f_{m,0}^{\alpha}$ and each $u_{i}$ has spectral density $f_{m,i}^{\alpha}$ for $i=1,2,\dotsc,m$. All these spectral densities are reciprocals of polynomials, implying that each process $u_{i}$, for $i=0,\ldots,m$, is a Gaussian Markov process (Pitt, 1971, Theorem 10.1). Specifically, $u_{0}$ is a Markov process of order $\max(\lfloor\alpha\rfloor,1)$, and it is $\max(\lfloor\alpha\rfloor-1,0)$ times differentiable in the mean-squared sense. Moreover, for $i>0$, each $u_{i}$ is a Markov process of order $\lceil\alpha\rceil$, and it is $\lfloor\alpha\rfloor$ times differentiable in the mean-squared sense. As a result, the multivariate process ${\boldsymbol{\mathrm{u}}}_{0}(t)=\left(u_{0}(t),u_{0}^{\prime}(t),\ldots,u_{0}^{(\max(\lfloor\alpha\rfloor-1,0))}(t)\right)$ is a first-order Markov process with a multivariate covariance function
${\boldsymbol{\mathrm{r}}}_{0}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{\lfloor\alpha\rfloor\times\lfloor\alpha\rfloor},\quad{\boldsymbol{\mathrm{r}}}_{0}(s,t)=\left[\frac{\partial^{i-1}}{\partial s^{i-1}}\frac{\partial^{j-1}}{\partial t^{j-1}}\varrho_{m,0}^{\alpha}(s-t)\right]_{i,j\in\{1,\ldots,\max(\lfloor\alpha\rfloor,1)\}}.$
@@
${\boldsymbol{\mathrm{r}}}_{i}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{\lceil\alpha\rceil\times\lceil\alpha\rceil},\quad{\boldsymbol{\mathrm{r}}}_{i}(s,t)=\left[\frac{\partial^{k-1}}{\partial s^{k-1}}\frac{\partial^{\ell-1}}{\partial t^{\ell-1}}\varrho_{m,i}^{\alpha}(s-t)\right]_{k,\ell\in\{1,\ldots,\lceil\alpha\rceil\}}.$
@@
Proposition 3.1.
Consider a set of locations $t_{1},\ldots,t_{n}\in I$. For $j=1,\ldots,n$, define the vectors
@@
${\boldsymbol{\mathrm{Q}}}_{i}=\begin{bmatrix}[l]{\boldsymbol{\mathrm{Q}}}_{1,1}&&{\boldsymbol{\mathrm{Q}}}_{1,2}&&&&&&\\ {\boldsymbol{\mathrm{Q}}}_{2,1}&&{\boldsymbol{\mathrm{Q}}}_{2,2}&&{\boldsymbol{\mathrm{Q}}}_{2,3}&&&&\\ &&{\boldsymbol{\mathrm{Q}}}_{3,2}&&{\boldsymbol{\mathrm{Q}}}_{3,3}&&{\boldsymbol{\mathrm{Q}}}_{3,4}&&\\ &&&\ddots&&\ddots&&\ddots&\\ &&&&{\boldsymbol{\mathrm{Q}}}_{n-1,n-2}&&{\boldsymbol{\mathrm{Q}}}_{n-1,n-1}&&{\boldsymbol{\mathrm{Q}}}_{n-1,n}\\ &&&&&&{\boldsymbol{\mathrm{Q}}}_{n,n-1}&&{\boldsymbol{\mathrm{Q}}}_{n,n}\end{bmatrix},$ (8)
for $i=0,1,\ldots,m$.
The first set of blocks, ${\boldsymbol{\mathrm{Q}}}_{1,1},{\boldsymbol{\mathrm{Q}}}_{1,2},$ and ${\boldsymbol{\mathrm{Q}}}_{2,1}$, are obtained as
@@
Remark 3.2.
If $\alpha\in\mathbb{N}$, there is no need for a rational approximation since the process itself is Markov. In this case, we can derive the precision matrix for the process and its derivatives using the same strategy as for ${\boldsymbol{\mathrm{u}}}_{i}$ in Proposition 3.1 but where the covariance function ${\boldsymbol{\mathrm{r}}}_{i}$ is replaced by the multivariate covariance function of $u$ and its derivatives. We do not go into more details in this case as there already are exact methods for this case. It should, however, be noted that the cost of using this exact method is the same as those we discuss below when setting $m=1$.
We now demonstrate how these expressions can be utilized for computationally efficient sampling, inference, and prediction. Let $t_{1},\ldots,t_{n}$ be a set of locations in $I$, and define the vector ${\boldsymbol{\mathrm{U}}}_{i}=\left[{\boldsymbol{\mathrm{u}}}_{i}(t_{1}),\ldots,{\boldsymbol{\mathrm{u}}}_{i}(t_{n})\right]$ for $i=0,\ldots,m$. Since these multivariate processes are independent, the vector $\bar{{\boldsymbol{\mathrm{U}}}}=\left[{\boldsymbol{\mathrm{U}}}_{0}^{\top},\ldots,{\boldsymbol{\mathrm{U}}}_{m}^{\top}\right]^{\top}$ is a centered multivariate Gaussian random variable of dimension $N=n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right),$ with a block diagonal precision matrix ${\boldsymbol{\mathrm{Q}}}=\operatorname{diag}({\boldsymbol{\mathrm{Q}}}_{0},{\boldsymbol{\mathrm{Q}}}_{1},\ldots,{\boldsymbol{\mathrm{Q}}}_{m})$, where each block is obtained using Proposition 3.1.
Since ${\boldsymbol{\mathrm{Q}}}$ is a band matrix, the following result provides the computational cost of computing its Cholesky factor.
Proposition 3.3.
@@
Partitioning $\bar{{\boldsymbol{\mathrm{U}}}}=(\bar{{\boldsymbol{\mathrm{U}}}}_{A}^{\top}\,\bar{{\boldsymbol{\mathrm{U}}}}_{B}^{\top})^{\top}$ for some partition $A\cup B=\{1,\ldots,N\}$ with $A\cap B=\emptyset$, we have that $\bar{{\boldsymbol{\mathrm{U}}}}_{A}|\bar{{\boldsymbol{\mathrm{U}}}}_{B}\sim\mathcal{N}(-{\boldsymbol{\mathrm{Q}}}_{A,A}^{-1}{\boldsymbol{\mathrm{Q}}}_{A,B}\bar{{\boldsymbol{\mathrm{U}}}}_{B},{\boldsymbol{\mathrm{Q}}}_{A,A}^{-1})$. This means that conditional distributions also can be computed efficiently, and in particular, the conditional mean ${\boldsymbol{\mathrm{\mu}}}_{A|B}=-{\boldsymbol{\mathrm{Q}}}_{A,A}^{-1}{\boldsymbol{\mathrm{Q}}}_{A,B}\bar{{\boldsymbol{\mathrm{U}}}}_{B}$ can be computed in $\mathcal{O}(|A|(2\lfloor\alpha\rfloor+1)^{2})$ computational cost.
@@
Proposition 3.4.
Using the same notation as in Proposition 3.1, the precision matrix of ${\boldsymbol{\mathrm{u}}}_{i}$ can be constructed as ${\boldsymbol{\mathrm{Q}}}_{i}={\boldsymbol{\mathrm{L}}}_{i}^{\top}{\boldsymbol{\mathrm{D}}}_{i}{\boldsymbol{\mathrm{L}}}_{i}$. Here ${\boldsymbol{\mathrm{D}}}_{i}$ is a diagonal matrix with positive diagonal entries and ${\boldsymbol{\mathrm{L}}}_{i}$ is a lower triangular block matrix with ones on the main diagonal. Specifically, ${\boldsymbol{\mathrm{L}}}_{i}$ has the form
@@
It should be noted that the costs $\mathcal{O}(n\max(\lfloor\alpha\rfloor,1)^{4})$ and $\mathcal{O}(n\lceil\alpha\rceil^{4})$ can likely be reduced slightly by taking advantage of that several redundant calculations are performed in the construction. Forming ${\boldsymbol{\mathrm{L}}}=diag({\boldsymbol{\mathrm{L}}}_{0},{\boldsymbol{\mathrm{L}}}_{1},\ldots,{\boldsymbol{\mathrm{L}}}_{m})$ and ${\boldsymbol{\mathrm{D}}}=diag({\boldsymbol{\mathrm{D}}}_{0},{\boldsymbol{\mathrm{D}}}_{1},\ldots,{\boldsymbol{\mathrm{D}}}_{m})$, where the blocks are obtained by using Proposition 3.4, we can now, for example, sample $\bar{{\boldsymbol{\mathrm{U}}}}$ as $\bar{{\boldsymbol{\mathrm{U}}}}={\boldsymbol{\mathrm{L}}}^{-1}{\boldsymbol{\mathrm{D}}}^{-\nicefrac{{1}}{{2}}}{\boldsymbol{\mathrm{Z}}}$ at $\mathcal{O}(nm\lceil\alpha\rceil^{2})$ computational cost.
Next, consider the situation of a Gaussian process regression where the stochastic process is observed under Gaussian measurement noise. That is, we have observations ${\boldsymbol{\mathrm{y}}}=[y_{1},y_{2},\dotsc,y_{n}]$ obtained as $y_{i}|u(\cdot)\sim\mathcal{N}(u(t_{i}),\sigma_{e}^{2}).$ Then, ${\boldsymbol{\mathrm{y}}}|\bar{{\boldsymbol{\mathrm{U}}}}\sim\mathcal{N}({\boldsymbol{\mathrm{A}}}\bar{{\boldsymbol{\mathrm{u}}}},\sigma_{e}^{2}\textbf{I})$ and ${{\boldsymbol{\mathrm{\bar{{\boldsymbol{\mathrm{U}}}}}}}\sim\mathcal{N}(0,{\boldsymbol{\mathrm{Q}}}^{-1})}$, where the matrix ${\boldsymbol{\mathrm{A}}}$ and the precision matrix ${\boldsymbol{\mathrm{Q}}}$ are same as defined above. The goal is typically to estimate the latent process by computing the posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}=\mathbb{E}({\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}})$ of the vector ${\boldsymbol{\mathrm{u}}}=(u(t_{1}),\ldots,u(t_{n}))^{\top}$.
@@
Computing this requires computing the Cholesky factors of ${\boldsymbol{\mathrm{Q}}}$ and ${\boldsymbol{\mathrm{Q}}}_{\bar{{\boldsymbol{\mathrm{U}}}}|{\boldsymbol{\mathrm{y}}}}$, after which we obtain the log-determinants and the posterior mean in linear cost. Therefore, the total cost for evaluating the log-likelihood, and thus, performing likelihood-based inference is $\mathcal{O}(nm^{3}\lceil\alpha\rceil^{3})$. To conclude, all relevant tasks for applying the proposed rational approximation in statistical inference require at most $\mathcal{O}(nm^{3}\lceil\alpha\rceil^{3})$ computational cost, and is thus linear in $n$.
Remark 3.6.
Because the rational approximation of the previous section results in a Gaussian process with a spectral density that is a rational function, an alternative to the Markov representations above is to directly apply the state space methods of (Karvonen & Särkkä, 2016) to perform inference and prediction at a linear cost $\mathcal{O}(M^{3}n)$, where $M$ is the approximation order for the state space method. However, the Markov approximation can directly be used in general Bayesian inference software such as R-INLA (Rue et al., 2009), which is not possible for the state space methods.

### v2

@@ 3 Linear cost inference
Let $\varrho_{m,\alpha}(\cdot)$ be defined as $\varrho_{m,\alpha}(t-s)=r_{m,\alpha}(t,s)$. Based on (8), we obtain the following explicit expression of the approximated covariance function $\varrho_{m,\alpha}(\cdot)$.
Proposition 5
@@
$\varrho_{m,\alpha}(h)=\varrho_{m,0,\alpha}(h)+\displaystyle\sum_{i=1}^{m}\varrho_{m,i,\alpha}(h),$ (9)
@@
$\varrho_{m,0,\alpha}(h)=k\sigma^{2}\cdot\begin{cases}\frac{c_{\alpha}\sqrt{4\pi}}{\kappa}1_{[h=0]}&0<\alpha<1,\\ \varrho\left(h;\lfloor\alpha\rfloor-\frac{1}{2},\kappa,\frac{c_{\alpha}}{c_{\lfloor\alpha\rfloor}}\right)&\alpha\geq 1,\end{cases}$
@@
$\varrho_{m,i,\alpha}(h)=c_{i}\sigma^{2}\cdot\begin{cases}\varrho\left(h;\frac{1}{2},\kappa_{i},\frac{c_{\alpha}\sqrt{\pi}}{\sqrt{1-p_{i}}}\right)&0<\alpha<1,\\ \frac{1}{p_{i}^{\lfloor\alpha\rfloor}}\varrho\left(h;\frac{1}{2},\kappa_{i},\frac{c_{\alpha}\sqrt{\pi}}{\sqrt{1-p_{i}}}\right)-\displaystyle\sum_{j=1}^{\lfloor\alpha\rfloor}\frac{1}{p_{i}^{\lfloor\alpha\rfloor+1-j}}\varrho\left(h;j-\frac{1}{2},\kappa,\frac{c_{\alpha}}{c_{j}}\right)&\alpha\geq 1.\end{cases}$
@@
Since the spectral density in (8) is a sum of valid spectral densities $f_{m,0,\alpha}$ and $f_{m,i,\alpha}$ for $i=1,\ldots,m$, a Gaussian process with the spectral density given by (8) can be expressed as a sum of independent Gaussian processes
@@
where $u_{0}$ has spectral density $f_{m,0,\alpha}$ and each $u_{i}$ has spectral density $f_{m,i,\alpha}$ for $i=1,2,\dotsc,m$. All these spectral densities are reciprocals of polynomials, implying that each process $u_{i}$, for $i=0,\ldots,m$, is a Gaussian Markov process (Pitt, 1971, Theorem 10.1). Specifically, $u_{0}$ is a Markov process of order $\max(\lfloor\alpha\rfloor,1)$, and it is $\max(\lfloor\alpha\rfloor-1,0)$ times differentiable in the mean-squared sense. Moreover, for $i>0$, each $u_{i}$ is a Markov process of order $\lceil\alpha\rceil$, and it is $\lfloor\alpha\rfloor$ times differentiable in the mean-squared sense. As a result, the multivariate process ${\boldsymbol{\mathrm{u}}}_{0}(t)=\left(u_{0}(t),u_{0}^{\prime}(t),\ldots,u_{0}^{(\max(\lfloor\alpha\rfloor-1,0))}(t)\right)$ is a first-order Markov process with a multivariate covariance function
${\boldsymbol{\mathrm{r}}}_{0}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{\lfloor\alpha\rfloor\times\lfloor\alpha\rfloor},\quad{\boldsymbol{\mathrm{r}}}_{0}(s,t)=\left[\frac{\partial^{i-1}}{\partial s^{i-1}}\frac{\partial^{j-1}}{\partial t^{j-1}}\varrho_{m,0,\alpha}(s-t)\right]_{i,j\in\{1,\ldots,\max(\lfloor\alpha\rfloor,1)\}}.$
@@
${\boldsymbol{\mathrm{r}}}_{i}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{\lceil\alpha\rceil\times\lceil\alpha\rceil},\quad{\boldsymbol{\mathrm{r}}}_{i}(s,t)=\left[\frac{\partial^{k-1}}{\partial s^{k-1}}\frac{\partial^{\ell-1}}{\partial t^{\ell-1}}\varrho_{m,i,\alpha}(s-t)\right]_{k,\ell\in\{1,\ldots,\lceil\alpha\rceil\}}.$
@@
Proposition 6
Consider a set of unique locations $t_{1},\ldots,t_{n}\in I$. For $j=1,\ldots,n$, define the vectors
@@
[…]
for $i=0,1,\ldots,m$. The first set of blocks, ${\boldsymbol{\mathrm{Q}}}_{1,1},{\boldsymbol{\mathrm{Q}}}_{1,2},$ and ${\boldsymbol{\mathrm{Q}}}_{2,1}$, are obtained as
@@
Remark 7
If $\alpha\in\mathbb{N}$, there is no need for a rational approximation since the process itself is Markov. In this case, we can derive the precision matrix for the process and its derivatives using the same strategy as for ${\boldsymbol{\mathrm{u}}}_{i}$ in Proposition 6 but where the covariance function ${\boldsymbol{\mathrm{r}}}_{i}$ is replaced by the multivariate covariance function of $u$ and its derivatives. We do not go into more details as there already are exact methods for $\alpha\in\mathbb{N}$. It should, however, be noted that the costs of using this exact method are the same as those we discuss below with $m=1$.
We now demonstrate how these expressions can be utilized for computationally efficient sampling, inference, and prediction. Let $t_{1},\ldots,t_{n}$ be a set of locations in $I$, and define the vector ${\boldsymbol{\mathrm{U}}}_{i}=\left[{\boldsymbol{\mathrm{u}}}_{i}(t_{1}),\ldots,{\boldsymbol{\mathrm{u}}}_{i}(t_{n})\right]$ for $i=0,\ldots,m$. Since these multivariate processes are independent, the vector $\bar{{\boldsymbol{\mathrm{U}}}}=\left[{\boldsymbol{\mathrm{U}}}_{0}^{\top},\ldots,{\boldsymbol{\mathrm{U}}}_{m}^{\top}\right]^{\top}$ is a centered multivariate Gaussian random variable of dimension $N=n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right),$ with a block diagonal precision matrix ${\boldsymbol{\mathrm{Q}}}=\operatorname{diag}({\boldsymbol{\mathrm{Q}}}_{0},{\boldsymbol{\mathrm{Q}}}_{1},\ldots,{\boldsymbol{\mathrm{Q}}}_{m})$, where each block is obtained using Proposition 6.
Because ${\boldsymbol{\mathrm{Q}}}$ is a band matrix, the following result provides the computational cost of computing its Cholesky factor.
Proposition 8
@@
Partitioning $\bar{{\boldsymbol{\mathrm{U}}}}=(\bar{{\boldsymbol{\mathrm{U}}}}_{A}^{\top}\,\bar{{\boldsymbol{\mathrm{U}}}}_{B}^{\top})^{\top}$ for some $A\cup B=\{1,\ldots,N\}$ with $A\cap B=\emptyset$, we have that $\bar{{\boldsymbol{\mathrm{U}}}}_{A}|\bar{{\boldsymbol{\mathrm{U}}}}_{B}\sim\mathcal{N}(-{\boldsymbol{\mathrm{Q}}}_{A,A}^{-1}{\boldsymbol{\mathrm{Q}}}_{A,B}\bar{{\boldsymbol{\mathrm{U}}}}_{B},{\boldsymbol{\mathrm{Q}}}_{A,A}^{-1})$. This means that conditional distributions also can be computed efficiently, and in particular, the conditional mean ${\boldsymbol{\mathrm{\mu}}}_{A|B}=-{\boldsymbol{\mathrm{Q}}}_{A,A}^{-1}{\boldsymbol{\mathrm{Q}}}_{A,B}\bar{{\boldsymbol{\mathrm{U}}}}_{B}$ can be computed in $\mathcal{O}(|A|(2\lfloor\alpha\rfloor+1)^{2})$ computational cost.
@@
Proposition 9
Using the same notation as in Proposition 6, the precision matrix of ${\boldsymbol{\mathrm{u}}}_{i}$ can be constructed as ${\boldsymbol{\mathrm{Q}}}_{i}={\boldsymbol{\mathrm{L}}}_{i}^{\top}{\boldsymbol{\mathrm{D}}}_{i}{\boldsymbol{\mathrm{L}}}_{i}$. Here ${\boldsymbol{\mathrm{D}}}_{i}$ is a diagonal matrix with positive diagonal entries and ${\boldsymbol{\mathrm{L}}}_{i}$ is a lower triangular block matrix with ones on the main diagonal. Specifically, ${\boldsymbol{\mathrm{L}}}_{i}$ has the form
@@
It should be noted that the costs $\mathcal{O}(n\max(\lfloor\alpha\rfloor,1)^{4})$ and $\mathcal{O}(n\lceil\alpha\rceil^{4})$ can likely be reduced slightly by taking advantage of that several redundant calculations are performed in the construction. Forming ${\boldsymbol{\mathrm{L}}}=diag({\boldsymbol{\mathrm{L}}}_{0},{\boldsymbol{\mathrm{L}}}_{1},\ldots,{\boldsymbol{\mathrm{L}}}_{m})$ and ${\boldsymbol{\mathrm{D}}}=diag({\boldsymbol{\mathrm{D}}}_{0},{\boldsymbol{\mathrm{D}}}_{1},\ldots,{\boldsymbol{\mathrm{D}}}_{m})$, where the blocks are obtained by using Proposition 9, we can now, for example, sample $\bar{{\boldsymbol{\mathrm{U}}}}$ as $\bar{{\boldsymbol{\mathrm{U}}}}={\boldsymbol{\mathrm{L}}}^{-1}{\boldsymbol{\mathrm{D}}}^{-\nicefrac{{1}}{{2}}}{\boldsymbol{\mathrm{Z}}}$ at $\mathcal{O}(nm\lceil\alpha\rceil^{2})$ computational cost.
Next, consider the situation of a Gaussian process regression where the stochastic process is observed under Gaussian measurement noise. That is, suppose that we have observations ${\boldsymbol{\mathrm{y}}}=[y_{1},y_{2},\dotsc,y_{n}]$ obtained as $y_{i}|u(\cdot)\sim\mathcal{N}(u(t_{i}),\sigma_{e}^{2}).$ Then, ${\boldsymbol{\mathrm{y}}}|\bar{{\boldsymbol{\mathrm{U}}}}\sim\mathcal{N}({\boldsymbol{\mathrm{A}}}\bar{{\boldsymbol{\mathrm{u}}}},\sigma_{e}^{2}\textbf{I})$ and ${{\boldsymbol{\mathrm{\bar{{\boldsymbol{\mathrm{U}}}}}}}\sim\mathcal{N}(0,{\boldsymbol{\mathrm{Q}}}^{-1})}$, where the matrix ${\boldsymbol{\mathrm{A}}}$ and the precision matrix ${\boldsymbol{\mathrm{Q}}}$ are same as defined above. The goal is typically to estimate the latent process by computing the posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}=\mathbb{E}({\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}})$ of the vector ${\boldsymbol{\mathrm{u}}}=(u(t_{1}),\ldots,u(t_{n}))^{\top}$.
@@
Computing this requires computing the Cholesky factors of ${\boldsymbol{\mathrm{Q}}}$ and ${\boldsymbol{\mathrm{Q}}}_{\bar{{\boldsymbol{\mathrm{U}}}}|{\boldsymbol{\mathrm{y}}}}$, after which we obtain the log-determinants and the posterior mean in linear cost. Therefore, the total cost for evaluating the log-likelihood, and thus, performing likelihood-based inference is $\mathcal{O}(nm^{3}\lceil\alpha\rceil^{3})$. To conclude, all relevant tasks for applying the proposed rational approximation in statistical inference require at most $\mathcal{O}(nm^{3}\lceil\alpha\rceil^{3})$ computational cost, and are thus linear in $n$.
Remark 11
Because the rational approximation of the previous section results in a Gaussian process with a spectral density that is a rational function, an alternative to the Markov representations above is to directly apply the state space methods of Karvonen and Särkkä (2016) to perform inference and prediction at a linear cost $\mathcal{O}(M^{3}n)$, where $M$ is the approximation order for the state space method. However, the Markov approximation can directly be used in general Bayesian inference software such as R-INLA (Rue et al., 2009), which is not possible for the state space methods.